OpenStudy (anonymous):

Which point is the solution to the system of equations? 8x + y = -16 -3x + y = -5 A. (-1, -8) B. (-2, 0) C. (1, -2) D. (-3, 8)

OpenStudy (anonymous):

Isolate y in both equation and equals them -16-8x=-5+3x 11x=-11 X=-1 Now substitute x i one of the equation Y=-16-8x Y=-16+8 Y=-8 Latter a is right